A cable lowers a 1400 kg elevator so that the elevator's speed increases from zero to 4.0 m/s in a vertical distance of 5.6 m .

What is the force that the cable exerts on the elevator while lowering it? (Assume that the system is the elevator and Earth.)

What is the force that the cable exerts on the elevator while lowering it? (Assume that the system is the elevator only.)
Answer #1

here,

initial speed , u = 0

final speed , v = 4 m/s

vertical distance , h = 5.6 m

let the accelration be a

v^2 - u^2 = 2 * a * h

4^2 - 0 = 2 * a * 5.6

a = 1.43 m/s^2

let the tension in the rope be t

m * g - t = m * a

1400 * 9.8 - t = 1400 * 1.43

t = 11718 N

the force that the cable exerts on the elevator while lowering it is 11718 N for both parts
Answer #2

A cable lowers a 1300 kg elevator so that the elevator's speed increases from zero to 4.0 m/s in a vertical distance of 6.4 m.


answered by: Madison Beck